What is the focus and directrix of the equation 2x^2 +15y=0?
Answer by lwsshak3(11628)

What is the focus and directrix of the equation
2x^2 +15y=0?
2x^2 =-15y
x^2=-15y/2
This is an equation of a parabola which opens downwards.
Its standard form: (x-h)^2=4p(y-k), (h,k)=(x,y) coordinates of the vertex
For given equation, x^2=-15y/2
vertex: (0,0)
axis of symmetry: y-axis or x=0
4p=15/2
p=15/8
focus: (0,-15/8) (p distance below vertex on axis of symmetry)
directrix: y=15/8 (p distance above vertex on axis of symmetry)